my son needs help with these also
7n+14-9n=26
5x=4=x-4

Answer by Bb(20)

ok for the 1st problem 7n+14-9n=26 u -14 on both sides and get 7n-9n=12 then u do simplify 7n-9n and u get -2n=12 then u divide both sides by -2 and u get your answer -6 to check your answer place -6 in for the n's and the problem will equal 26 which is the answer u want



for the 2nd problem 5x=4=x-4 u +4 to the 4 and u get 5x=8=x u subtract the x from 5x and get 4x=8 u divde both sides by 4 and u get x=2 i'm not exactly positive on on this one but i'm pretty sure that's how u do it



i hope that helped u